NicktheBlade16:

PLZ HELP, IT'S ALGEBRA 2:Which of the following is the solution to the equation 25(z − 4) = 125? z = 5.5 z = 3.5 z = −2.5 z = −0.5

Alexis1415:

The answer would be A, z=5.5 Reasoning: We have our equation '25(z − 4) = 125" So, \[5^{2} (z-4)=5^{3}\] \[2(z-4) =3\] \[3+8=11 \] So 2z would equal 11 11 divided by 2 gives us 5.5 I hope this helps c:

NicktheBlade16:

Wow ok, I can get that thank you so much. Btw, i might have more

Alexis1415:

@nicktheblade16 wrote:
Wow ok, I can get that thank you so much. Btw, i might have more
ofccc, and yeah if you post anymore just @ me